# SimBench: Benchmarking the Ability of Large Language Models to Simulate Human Behaviors

### v3

@@ 4.2 RQ2: Impact of LLM Characteristics on Simulation Ability
While even the best models struggle to perform well on SimBench, Table 1 also shows clear differences across models. Therefore, we investigate how performance varies depending on model characteristics, specifically 1) model size, and 2) test-time compute.
@@
To evaluate the impact of model size on simulation ability, we plot SimBench Score $S$ against model parameter count for the four LLM families that we can test across multiple model sizes (Figure 2 and Appendix Figure 6). Our results suggest that there is a clear log-linear scaling law for LLM simulation ability. Across all examined model families, for both base and instruction models, an increase in parameter count generally corresponds to an increase in SimBench score $S$, indicating better alignment between predicted and human response distributions. There is also an interaction between model size and the effect of instruction-tuning. While instruction-tuned models consistently outperform their base counterparts at larger scales (>10B parameters), this relationship appears to invert for smaller models. For example, the OLMo-2 base models outperform their instruction-tuned variants at the 7B and 13B scale. Furthermore, the plot shows that instruction-tuned models not only reach a higher peak performance but also appear to scale more effectively. The steeper slope of the dashed lines (e.g., for Qwen2.5-Instruct) compared to the solid lines suggests that instruction-tuning may improve a model’s ability to capitalize on increases in parameter count for the simulation task. We present a more comprehensive plot including all evaluated model families in Appendix 6, which confirms this trend holds across considered models.
Overall, the clear positive scaling trends across model families suggest that, while simulation remains a challenging task for even the best models today, further model scaling may well lead to highly accurate LLM simulators in the future.
2) Test-Time Compute
@@
To analyze the effects of increasing test-time compute on LLM simulation ability, we conduct two sets of experiments. We compare the performance of two distinct o4-mini checkpoints (‘low‘ vs. ‘high‘, which vary in the amount of reasoning efforts), and we assess Claude-3.7-Sonnet with and without a 4000-token reasoning budget. Additionally, we apply a zero-shot Chain-of-Thought (CoT) prompting strategy (Wei et al., 2022) to GPT-4.1 and DeepSeek-V3-0324 (see Appendix D for prompt details).
@@
Our results suggest that increasing test-time compute provides no meaningful benefit for LLM simulation ability. The o4-mini model has a minor improvement (S score: 28.20 → 29.54), while Claude-3.7-Sonnet’s performance slightly decreases (40.51 → 39.46). Similarly, applying CoT prompting leads to a small performance drop for GPT-4.1 (34.90 → 33.11) and a negligible change for DeepSeek-V3-0324 (33.14 → 33.16). This result aligns with a growing body of recent work showing that the benefits of test-time compute are highly task-dependent (Liu et al., 2024; Sprague et al., 2025; Gema et al., 2025) and do not necessarily improve role-playing ability (Feng et al., 2025). This may be because CoT enforces a mode of explicit, step-by-step reasoning that is well-documented to harm performance on tasks that rely on faster, more intuitive human heuristics Liu et al. (2024).
@@ 4.3 RQ3: Impact of Task Selection on Simulation Fidelity
The 20 datasets in SimBench correspond to very different tasks, in terms of the aspects of human behavior that they measure (see §2.1). Therefore, we break down simulation fidelity by dataset, showing results for the five LLMs we previously identified as the best simulators in Figure 3. We find that simulation fidelity varies substantially across tasks, with even the best LLM simulators performing worse than a uniform response baseline on several datasets, as indicated by negative SimBench scores (e.g., on Jester, OSPsychMach, and MoralMachine). Generally, most LLMs exhibit similar performance patterns, though GPT-4.1 scores exceptionally high (61.9) on OSPsychRWAS.
@@ 4.4 RQ4: The Alignment-Simulation Tradeoff
Faithful simulation requires models to capture the full spectrum of human opinion, from strong consensus to widespread disagreement. We operationalize this “response plurality” using the normalized entropy of the human response distribution. Prior work suggests that standard alignment via instruction-tuning encourages confident, low-entropy outputs (Brown et al., 2020; Tian et al., 2023; Meister et al., 2025; Cruz et al., 2024; Hu et al., 2025b; Dong et al., 2025), creating a potential conflict with simulating diverse human perspectives.
An Systematic Tradeoff between Alignment and Plurality.
@@
Our analysis confirms this hypothesis, revealing a systematic tradeoff. As detailed in Appendix I, base models consistently outperform their instruction-tuned counterparts on high-entropy questions, while the inverse is true for low-entropy questions. To precisely quantify this effect, we compute the change in SimBench score ($\Delta S=S_{\text{instruct}}-S_{\text{base}}$) for 13 model pairs. Figure 4 visualizes the resulting trend by aggregating all data points into 25 bins based on human response entropy and plotting the average improvement for each. The result is a near-perfect negative linear relationship (r=0.942r=-0.942), revealing two distinct regimes. On low-entropy questions where humans agree, post-training provides a substantial benefit, improving the S-score by up to 40 points. This is alignment working as intended. However, as human disagreement (entropy) increases, the benefit of instruction-tuning systematically erodes, crossing a point of no-improvement around an entropy of 0.8. For questions with the highest plurality, post-training becomes detrimental, making the aligned model a worse simulator than its base counterpart.
@@
This empirical finding is well-explained by the theoretical framework of reinforcement learning (RL) as Bayesian inference (Levine, 2018; Korbak et al., 2022) The pre-training objective of a base model typically minimizes a mass-covering KL divergence ($D_{KL}(p\|q)$), which encourages the model ($q$) to place probability mass wherever the true data distribution ($p$) has mass. This process naturally leads to models that represent the full, multi-modal diversity of human language and opinion seen in their training data. In contrast, alignment via KL-regularized RL (e.g., RLHF) minimizes a mode-seeking KL divergence ($D_{KL}(q\|\sigma)$). This objective incentivizes the model ($q$) to find and concentrate its probability mass on a single, high-reward mode of the target preference distribution ($\sigma$), even at the cost of ignoring other valid modes. Our results provide strong empirical validation of this theoretical distinction: alignment optimizes for a single “best” response, fundamentally training the model to discard the pluralistic, high-entropy distributions characteristic of genuine human populations.
Decomposing Instruction Tuning’s Dual Effects: Helpful Instruction Following vs. Harmful Entropy Reduction
@@
To formally test this mechanism, a causal mediation analysis (Appendix I.3) decomposes the effect of instruction-tuning into two larger, opposing forces: a large, positive direct effect on performance (+6.46), likely from improved instruction-following, and a significant, negative indirect effect (-1.74) mediated by the model’s reduced output entropy.
@@
This tradeoff between instruction-following and diversity raises a critical question: are there other ways to improve simulation?To investigate this, we contrast general-purpose alignment with the specialist cognitive tuning of the Centaur models (Binz et al., 2025). Centaur models are Llama models fine-tuned on Psych-101, a large dataset of lab experiments, making our diverse SimBench a powerful out-of-distribution test of their generalization. Both approaches improve simulation over the Llama-3.1-70B base model, but they do so via opposing mechanisms (Section K). General-purpose instruction tuning ($S=16.56$) leverages the helpful direct effect of alignment, excelling on low-entropy consensus tasks. In contrast, specialist cognitive tuning ($S=8.54$) improves performance by avoiding the harmful indirect effect, preserving the base model’s intrinsic ability to capture high-entropy pluralistic responses. The existence of these two distinct—and currently separate—paths to improving simulation underscores a key challenge and opportunity: the most faithful simulators of the future will likely need to synthesize the benefits of both general-purpose alignment and distribution-preserving cognitive modeling.
Figure 4: The alignment-simulation tradeoff: Instruction-tuning helps on consensus questions but hurts on diverse ones. The plot aggregates results from 13 base/instruction-tuned model pairs. Each point represents the average improvement in SimBench Score for one of 25 entropy bins. The point’s x-coordinate is the mean entropy of all questions within that bin. Error bars show the standard error of the mean.
@@ 4.5 RQ5: Simulation Ability Across Participant Groups
Table 2: Ungrouped vs. grouped simulation performance ($\Delta S$).
@@
Claude-3.7-Sonnet -3.13
Claude-3.7-Sonnet-4000 -4.61
DeepSeek-R1 -3.79
DeepSeek-V3-0324 -1.27
GPT-4.1 -3.94
@@
Religiosity/Practice -9.91
Political Affil./Ideology -4.97
Religion (Affiliation) -4.83
Income/Social Standing -4.51
Domicile/Urbanicity -3.17
Employment Status -3.03
Education -2.55
Marital Status -1.80
Age -1.50
Gender -1.24
We measure this change as $\Delta S=S_{grouped}-S_{ungrouped}$, where $S_{ungrouped}$ is the SimBench score for simulating the general population and $S_{grouped}$ is the score when simulating a specific demographic group on the same question. A negative $\Delta S$ indicates that the model’s simulation ability relative to the uniform baseline decreases when asked to simulate specific demographic groups.
Importantly, for SimBenchGrouped, we specifically selected questions where human response distributions showed the highest variance across demographic groups (see §2.4). The observed degradation in simulation performance therefore likely represents an upper bound on the challenges LLMs face when simulating specific demographic groups. Our results in Table 2 show that LLMs struggle more with simulating specific demographic groups compared to general populations. All evaluated models show negative mean $\Delta S$ values, with degradation ranging from -1.27 for DeepSeek-V3-0324 to -4.61 for Claude-3.7-Sonnet-4000.
@@
Finally, we analyze the relationship between LLM simulation ability and more general model capabilities by correlating performance on SimBench with popular LLM capability benchmarks. We collect performance data for eight models on five benchmarks representing distinct capabilities and calculate the Pearson correlation with their SimBench scores (see Appendix J for detailed implementation details and scatter plots).
We find that SimBench performance correlates most strongly with benchmarks requiring deep and broad knowledge-intensive reasoning, such as MMLU-Pro (r = 0.94) and GPQA Diamond (r = 0.86). This relationship is weaker for chat helpfulness as measured by Chatbot Arena ELO (r = 0.71) and instruction following (IF-Eval, r = 0.79). Crucially, the correlation is substantially weaker for narrow, specialized skills like advanced mathematics (OTIS AIME, r = 0.48). These results indicate that accurately simulating human behavior is a complex capability rooted in broad, knowledge-intensive reasoning, which aligns with the diverse social and behavioral topics in SimBench. The weaker correlations with Chatbot Arena and AIME show that neither general chat ability nor narrow problem-solving skills are sufficient proxies for strong simulation performance.
@@ 5 Related Work
LLMs as human behavior simulators have attracted significant interdisciplinary attention. Researchers have evaluated their efficacy across political science (Argyle et al., 2023; Bisbee et al., 2024; Dominguez-Olmedo et al., 2024), psychology (Aher et al., 2023; Binz et al., 2025; Manning et al., 2024; Hewitt et al., 2024; Rescala et al., 2024), economics (Horton, 2023; Aher et al., 2023), and computer science applications (Hu & Collier, 2024; Dong et al., 2024; Hu & Collier, 2025; Park et al., 2023). Evidence regarding LLMs’ simulation fidelity remains mixed, with some studies reporting promising results (Argyle et al., 2023) while others identify critical limitations, including homogenized group representations (Cheng et al., 2023; Wang et al., 2025) and deterministic rather than distributional predictions (Park et al., 2024b).
Existing work has predominantly focused on individual-level simulation with minimal demographic conditioning, typically evaluating only one or two models in narrowly defined contexts. SimBench addresses these limitations by providing a comprehensive benchmark for group-level simulation across diverse domains with systematic demographic conditioning and standardized metrics. The benchmark’s distributional evaluation framework (using Total Variation distance) captures how accurately models represent the full spectrum of human response variation – an approach advocated by researchers in both simulation (Anthis et al., 2025) and general LLM evaluation (Ying et al., 2025). For broader context on this emerging field, we refer readers to recent comprehensive surveys (Kozlowski & Evans, 2024; Olteanu et al., 2025; Anthis et al., 2025).

### v4

@@ 4.2 RQ2: Impact of LLM Characteristics on Simulation Ability
While even the best models struggle to perform well on SimBench, Table 1 also shows clear differences across models. Therefore, we investigate how performance varies depending on model characteristics, specifically 1) model size, and 2) inference-time compute.
@@
To evaluate the impact of model size on simulation ability, we plot SimBench Score $S$ against model parameter count for the four LLM families that we can test across multiple model sizes (Figure 2 and Appendix Figure 6). Our results suggest that there is a log-linear scaling trend for LLM simulation ability. Across examined model families, for both base and instruction models, an increase in parameter count generally corresponds to an increase in SimBench score $S$, indicating better alignment between predicted and human response distributions. This relationship is robustly supported by model families with comprehensive size variants (e.g., Qwen2.5, Llama-3.1), though additional data points would be needed to fully characterize the trajectory for families with fewer models (e.g., OLMo). There is also an interaction between model size and the effect of instruction tuning. While instruction-tuned models consistently outperform their base counterparts at larger scales (>10B parameters), this relationship appears to invert for smaller models. For example, the OLMo-2 base models outperform their instruction-tuned variants at the 7B and 13B scale. Furthermore, the plot shows that instruction-tuned models not only reach a higher peak performance but also appear to scale more effectively. The steeper slope of the dashed lines (e.g., for Qwen2.5-Instruct) compared to the solid lines suggests that instruction tuning may improve a model’s ability to capitalize on increases in parameter count for the simulation task. We present a more comprehensive plot including all evaluated model families in Appendix G, which confirms this trend holds across models.
Overall, the clear positive scaling trends across model families suggest that simulation ability is a capability that improves with scale. However, the log-linear nature of this relationship implies diminishing returns from scale alone, reinforcing the need for targeted approaches — such as distribution-preserving alignment (§4.4) — to achieve substantial further gains.
Figure 2: Model parameter count vs. simulation ability. We measure model size by parameter count and simulation ability by SimBench score $S$ averaged across the two main splits of SimBench. Shaded bands denote 95% confidence intervals around the mean score.
2) Inference-Time Compute
@@
To analyze the effects of increasing inference-time compute on LLM simulation ability, we conduct two sets of experiments. First, we compare the performance of two distinct o4-mini checkpoints (“low” vs. “high”, which vary in the amount of reasoning effort), as well as Claude-3.7-Sonnet without reasoning and with a 4000-token reasoning budget. Second, we apply a zero-shot Chain-of-Thought (CoT) prompting strategy (Wei et al., 2022) to GPT-4.1 and DeepSeek-V3-0324 (see Appendix D for prompt details).
@@
Our results suggest that, across the models tested, increasing inference-time compute provides no meaningful benefit for LLM simulation ability. The o4-mini model shows minor improvement (SimBench $S$ score: 27.77 → 28.99), while Claude-3.7-Sonnet’s performance slightly decreases (40.80 → 39.46). Similarly, applying CoT prompting leads to a small performance drop for GPT-4.1 (34.55 → 33.11) and a negligible change for DeepSeek-V3-0324 (32.89 → 33.16). While our analysis covers a limited number of models, this result aligns with a growing body of recent work showing that inference-time compute benefits are highly task-dependent (Liu et al., 2025b; Sprague et al., 2025; Gema et al., 2025) and do not necessarily improve role-playing ability (Feng et al., 2025). We hypothesize this is because CoT forces an overly rational deliberation that mismatches the often heuristic-driven nature of human responses in SimBench.
@@ 4.3 RQ3: Impact of Task Selection on Simulation Fidelity
The 20 datasets in SimBench correspond to tasks that are highly diverse in terms of which aspects of human behavior they measure (see §2.1). Breaking down simulation fidelity by dataset (Figure 3) reveals that simulation fidelity varies substantially across tasks. Models are most successful at simulating responses to standard survey questions regarding stated opinions, attitudes, and self-assessments (e.g., OpinionQA, Afrobarometer). However, performance degrades on tasks requiring the simulation of a behavioral choice, whether in risky choice problems (Choices13k) or moral dilemmas (MoralMachine). This finding provides large-scale evidence for a “value-action gap” in LLMs, echoing recent work (Shen et al., 2025) which suggests a misalignment exists between LLM-generated value statements and their actions.
Finally, models struggle severely to capture human response distributions on datasets involving traits or beliefs that conflict with standard alignment objectives. For even the best LLMs, we observe extremely poor performance, often worse than a uniform baseline, on datasets measuring Machiavellianism (OSPsychMACH), conspiratorial beliefs (ConspiracyCorr), or humor rating (Jester). This aligns with a growing body of work (Liu et al., 2024; Kumar et al., 2025; Yi et al., 2025) showing that alignment filters may inhibit the simulation of “atypical” or counter-normative human perspectives. While most LLMs exhibit these performance patterns, GPT-4.1 is a notable outlier, scoring exceptionally high (61.9) on OSPsychRWAS.
@@ 4.4 RQ4: The Alignment-Simulation Tradeoff
Faithful simulation requires models to capture the full spectrum of human opinion, from strong consensus to widespread disagreement. We operationalize this “response plurality” using the normalized entropy of the human response distribution.
An Empirical Tradeoff between Alignment and Plurality.
Prior work suggests that standard alignment via instruction tuning encourages confident, low-entropy outputs (Brown et al., 2020; Tian et al., 2023; Cruz et al., 2024; Meister et al., 2025), which creates a potential conflict with simulating diverse human perspectives. Our analysis reveals exactly this tradeoff. As detailed in Appendix I, base models consistently outperform their instruction-tuned counterparts on high-entropy questions, while the inverse is true for low-entropy questions. To precisely quantify this effect, we compute the change in SimBench score ($\Delta S=S_{\text{instruct}}-S_{\text{base}}$) for 13 model pairs and plot this performance gain from post-training against human response entropy.
@@
The result, shown in Figure 4, is a near-perfect negative linear relationship (r=0.942r=-0.942). This plot reveals two distinct regimes. On low-entropy questions where humans agree, post-training provides a substantial benefit, improving the S-score by up to 40 points. However, as human disagreement (entropy) increases, the benefit of instruction tuning systematically erodes, crossing a point of no-improvement around an entropy of 0.8. For questions with the highest plurality, post-training becomes actively detrimental, making the aligned model a worse simulator than its base counterpart.
@@
This empirical finding is well-explained by the theoretical framework of reinforcement learning (RL) as Bayesian inference (Levine, 2018; Korbak et al., 2022). The pre-training objective of a base model typically minimizes a mass-covering KL divergence ($D_{KL}(p\|q)$), which encourages the model ($q$) to place probability mass wherever the true data distribution ($p$) has mass. This process naturally leads to models that represent the full, multi-modal diversity of human language and opinion seen in their training data. In contrast, alignment via KL-regularized RL (e.g., RLHF) minimizes a mode-seeking KL divergence ($D_{KL}(q\|\sigma)$). This objective incentivizes the model ($q$) to find and concentrate its probability mass on a single, high-reward mode of the target preference distribution ($\sigma$), even at the cost of ignoring other valid modes. Our results provide strong empirical validation of this theoretical distinction: alignment optimizes for a single “best” response, effectively encouraging the model to discard the pluralistic, high-entropy distributions characteristic of genuine human populations.
@@
To formally test this mechanism, a causal mediation analysis (Appendix I.3) decomposes the effect of instruction tuning into two larger, opposing forces: a large, positive direct effect on performance (+6.46), likely from improved instruction-following, and a significant, negative indirect effect (-1.74) mediated by the model’s reduced output entropy.
@@
This formally-decomposed tradeoff is perfectly illustrated by comparing general-purpose instruction tuning with the specialist cognitive tuning of the Centaur models (Binz et al., 2025). Centaur models are Llama models fine-tuned on Psych-101, a large dataset of lab experiments, making our diverse SimBench a powerful out-of-distribution test of their generalization. Both approaches improve simulation over the Llama-3.1-70B base model, but they do so via opposing mechanisms. General-purpose instruction tuning ($S=16.56$) leverages the helpful direct effect of alignment, excelling on low-entropy consensus tasks. In contrast, specialist cognitive tuning ($S=8.54$) improves performance by avoiding the harmful indirect effect, preserving the base model’s intrinsic ability to capture high-entropy pluralistic responses. The existence of these two distinct—and currently separate—paths to improving simulation underscores a key challenge and opportunity: the most faithful simulators of the future will likely need to synthesize the benefits of both general-purpose alignment and distribution-preserving cognitive modeling. Other promising directions include post-hoc weight interpolation (Hu et al., 2025) and inference-time methods to amplify system prompt adherence beyond the default assistant persona (Dong et al., 2026).
Figure 4: The Alignment-Simulation Tradeoff: Instruction Tuning Helps on Consensus Questions but Hurts on Diverse ones. We bin questions by the normalized entropy of their human response distribution (x-axis) and plot the mean improvement in SimBench Score when switching from a base model to its instruction-tuned counterpart (y-axis).
@@ 4.5 RQ5: Simulation Ability Across Participant Groups
Table 2: Ungrouped vs. grouped simulation performance ($\Delta S$). $\pm$ denotes 95% CI.
@@
Claude-3.7-Sonnet ${-3.13}_{{\pm 0.66}}$
Claude-3.7-Sonnet-4000 ${-4.61}_{{\pm 0.65}}$
DeepSeek-R1 ${-3.79}_{{\pm 0.76}}$
DeepSeek-V3-0324 ${-1.27}_{{\pm 0.69}}$
GPT-4.1 ${-3.94}_{{\pm 0.76}}$
@@
Religiosity/Practice ${-9.91}_{{\pm 1.59}}$
Political Affil./Ideology ${-4.97}_{{\pm 1.28}}$
Religion (Affiliation) ${-4.83}_{{\pm 1.16}}$
Income/Social Standing ${-4.51}_{{\pm 1.23}}$
Domicile/Urbanicity ${-3.17}_{{\pm 0.82}}$
Employment Status ${-3.03}_{{\pm 0.82}}$
Education ${-2.55}_{{\pm 0.99}}$
Marital Status ${-1.80}_{{\pm 0.87}}$
Age ${-1.50}_{{\pm 0.85}}$
Gender ${-1.24}_{{\pm 0.72}}$
We measure this change as $\Delta S=S_{grouped}-S_{ungrouped}$, where $S_{ungrouped}$ is the SimBench score for simulating the general population and $S_{grouped}$ is the score when simulating a specific demographic group on the same question. A negative $\Delta S$ indicates that the model’s simulation ability decreases relative to the uniform baseline when it is asked to simulate specific demographic groups.
Importantly, for SimBenchGrouped, we specifically selected questions where human response distributions showed the highest variance across demographic groups (see §2.4). The observed degradation in simulation performance therefore likely represents an upper bound on the challenges LLMs face when simulating specific demographic groups. Our results in Table 2 show that LLMs struggle more with simulating specific demographic groups compared to general populations. All evaluated models show negative mean $\Delta S$ values, with degradation ranging from -1.27 for DeepSeek-V3-0324 to -4.61 for Claude-3.7-Sonnet-4000. All degradations are statistically significant, with 95% confidence intervals excluding zero in every case.
@@
Finally, we analyze the relationship between LLM simulation ability and more general model capabilities by correlating performance on SimBench with popular LLM capability benchmarks. We collect performance data for eight models on five benchmarks representing distinct capabilities and calculate the Pearson correlation with their SimBench scores (see Appendix J for implementation details and scatter plots).
We find that SimBench performance correlates most strongly with benchmarks requiring knowledge-intensive reasoning, such as MMLU-Pro (r = 0.94) and GPQA Diamond (r = 0.86). The correlation is weaker for general helpfulness (Chatbot Arena ELO, r = 0.71) and instruction following (IF-Eval, r = 0.79). Crucially, the correlation is substantially weaker for narrow, specialized skills like advanced mathematics (OTIS AIME, r = 0.48). We posit that accurately simulating human behavior is a complex capability rooted in broad, knowledge-intensive reasoning, which aligns with the diverse social and behavioral topics in SimBench. The weaker correlations with Chatbot Arena and OTIS AIME suggest that neither general conversational ability nor narrow problem-solving skills are sufficient proxies for strong simulation performance.
@@ 5 Related Work
LLMs as human behavior simulators have attracted significant interdisciplinary attention. Researchers have evaluated their efficacy across political science (Argyle et al., 2023; Bisbee et al., 2024; Dominguez-Olmedo et al., 2024), psychology (Aher et al., 2023; Manning et al., 2024; Hewitt et al., 2024; Binz et al., 2025), economics (Horton, 2023; Aher et al., 2023), and computer science applications (Park et al., 2023; Hu & Collier, 2024; Dong et al., 2024; Hu & Collier, 2025). Evidence regarding LLMs’ simulation fidelity remains mixed, with some studies reporting promising results (Argyle et al., 2023) while others identify critical limitations, including homogenized group representations (Cheng et al., 2023; Wang et al., 2025), a tendency toward hyper-rationality rather than human-like error (Liu et al., 2025a), and deterministic rather than distributional predictions (Park et al., 2024b).
Existing work has predominantly focused on individual-level simulation with minimal demographic conditioning, typically evaluating only one or two models in narrowly defined contexts. SimBench addresses these limitations by providing a comprehensive benchmark for group-level simulation across diverse domains with systematic demographic conditioning and standardized metrics. SimBench’s distributional evaluation framework, based on Total Variation Distance, captures how accurately models represent the full spectrum of human response variation. This is an approach advocated for by researchers in both simulation (Anthis et al., 2025) and general LLM evaluation studies (Ying et al., 2025). For broader context on this emerging field, we refer readers to recent comprehensive surveys (Ma et al., 2024; Kozlowski & Evans, 2025; Olteanu et al., 2025; Anthis et al., 2025).